# The Mean-Field Dynamics of Transformers

### v2

@@ Abstract.
Normalization layers are a defining component of modern Transformers. The original formulation of attention as an interacting particle system without normalization was introduced in [sander2022sinkformers]. In that setting tokens move freely in $\mathbb{R}^{d}$ and typically either diverge or collapse to the origin, and clustering results require additional assumptions and can be elicited on rescaled tokens [geshkovski2024emergence, castin2025unified]. By contrast, the presence of layer normalization (LN) as in real architectures keeps representations on a controlled scale and produces crisp clustering behavior.
@@
while the magnitudes $r_{i}(t)$ satisfy a rule-dependent radial equation. Crucially, each normalization scheme induces a corresponding speed regulation factor $s_{i}(t)$ on token $i$. For example, for Post-LN, taking $s_{i}(t)=1$ recovers (SA) while it can be shown that Pre-LN yields $s_{i}(t)=r_{i}(t)$ so that the directions of tokens with large magnitude are effectively slowed down. Other variants include Peri-LN [kim2025peri], Mix-LN [li2024mixln], nGPT [loshchilov2024ngpt], and sqrt-scaling [Nocisqrt]. All fit into this framework with different choices of $s_{i}(t)$.
Using this unified framework, [karagodin2025normalization] show that attention dynamics collapse to a single cluster in long time even in the presence of speed regulation factors. Moreover, studying these dynamics under the equiangular model yields a sharper image that reveals important distinctions between the various LN schemes.
@@
Figure 4. Evolution of average cosine similarity $\rho(t)$ with 90% confidence interval on a Transformer with random weights, $d=512,\beta=\sqrt{d},d>n$ and random initial tokens. Figure reproduced from [karagodin2025normalization].
The equiangular model appears to be predictive of the rate of clustering despite its apparent simplicity. In fact, the linearization argument may be applied beyond the equiangular setup and similar conclusions can be made in these more general dynamics [karagodin2025normalization, Theorem 4.3].
@@
The equiangular model also sheds light on the behavior of long-context Transformers, where $n$ is large and attention scores tend to flatten. In such regimes, the softmax denominator grows proportionally to $n$, and unless the attention scale grows with $n$, the weights $A_{ij}$ approach $1/n$ and drive the system toward uniform mixing as in the Kuramoto model. This phenomenon amplifies the contractive dynamics described above and accelerates clustering and hence representation collapse. Motivated by this effect, several practical long-context systems including Qwen [bai2023qwen], SSMax [nakanishi2025scalable], and SWAN-GPT [puvvada2025swan] adopt a logarithmic attention scaling $\beta_{n}=\Theta(\log n)$.
@@
The bifurcation structure of stationary solutions to this noisy Transformer model was first analyzed in [shalova2024solutions], extending earlier work in [carrillo2020long]. For $d=2$, these results were sharpened in [banerjee]. As in the deterministic setting, the noisy Transformer dynamics interpolate with the classical noisy Kuramoto model, obtained by taking $\beta=0$. This special case is known under various names—including the mean-field plane rotator and XY-spin models—and its stationary and dynamical properties are well understood: the uniform distribution is the unique stationary solution for $\kappa\leq 2$, a pitchfork bifurcation occurs at $\kappa=2$, and a unique (up to rotation) nontrivial branch exists for $\kappa>2$ [bertini2010dynamical]. A significant body of work, culminating in the uniform-in-time propagation-of-chaos result of [DelTse25], now gives a remarkably complete picture of the noisy Kuramoto model.
By contrast, a quantitative description of the noisy Transformer dynamics remains largely open. Phenomena such as metastability suggest that the strong uniform-in-time results available for the noisy Kuramoto model may not hold in this richer setting; see, e.g., [GarPapYan17, hairer]. Moreover, many natural variations, such as introducing common noise, exploring different geometries, modifying the interaction kernel, or considering anisotropic or multiplicative noise, lead to further mathematical challenges. Understanding these variants, and charting the full phase diagram of noisy attention dynamics, presents a wide landscape of open problems.
@@ Acknowledgments.
I am most grateful to Borjan Geshkovski, Shi Chen, and Zhengjiang Lin for many insightful comments on early versions of this text.

### v3

@@ Abstract.
Normalization layers are a defining component of modern Transformers. The original formulation of attention as an interacting particle system without normalization was introduced in [60]. In that setting tokens move freely in $\mathbb{R}^{d}$ and typically either diverge or collapse to the origin, and clustering results require additional assumptions and can be elicited on rescaled tokens [31, 14]. By contrast, the presence of layer normalization (LN) as in real architectures keeps representations on a controlled scale and produces crisp clustering behavior.
@@
while the magnitudes $r_{i}(t)$ satisfy a rule-dependent radial equation. Crucially, each normalization scheme induces a corresponding speed regulation factor $s_{i}(t)$ on token $i$. For example, for Post-LN, taking $s_{i}(t)=1$ recovers (SA) while it can be shown that Pre-LN yields $s_{i}(t)=r_{i}(t)$ so that the directions of tokens with large magnitude are effectively slowed down. Other variants include Peri-LN [41], Mix-LN [45], nGPT [50], and sqrt-scaling [54]. All fit into this framework with different choices of $s_{i}(t)$.
Using this unified framework, [39] show that attention dynamics collapse to a single cluster in long time even in the presence of speed regulation factors. Moreover, studying these dynamics under the equiangular model yields a sharper image that reveals important distinctions between the various LN schemes.
@@
Figure 4. Evolution of average cosine similarity $\rho(t)$ with 90% confidence interval on a Transformer with random weights, $d=512,\beta=\sqrt{d},d>n$ and random initial tokens. Figure reproduced from [39].
The equiangular model appears to be predictive of the rate of clustering despite its apparent simplicity. In fact, the linearization argument may be applied beyond the equiangular setup and similar conclusions can be made in these more general dynamics [39, Theorem 4.3].
@@
The equiangular model also sheds light on the behavior of long-context Transformers, where $n$ is large and attention scores tend to flatten. In such regimes, the softmax denominator grows proportionally to $n$, and unless the attention scale grows with $n$, the weights $A_{ij}$ approach $1/n$ and drive the system toward uniform mixing as in the Kuramoto model. This phenomenon amplifies the contractive dynamics described above and accelerates clustering and hence representation collapse. Motivated by this effect, several practical long-context systems including Qwen [6], SSMax [53], and SWAN-GPT [57] adopt a logarithmic attention scaling $\beta_{n}=\Theta(\log n)$.
@@
The bifurcation structure of stationary solutions to this noisy Transformer model was first analyzed in [61], extending earlier work in [13]. For $d=2$, these results were sharpened in [7]. As in the deterministic setting, the noisy Transformer dynamics interpolate with the classical noisy Kuramoto model, obtained by taking $\beta=0$. This special case is known under various names—including the mean-field plane rotator and XY-spin models—and its stationary and dynamical properties are well understood: the uniform distribution is the unique stationary solution for $\kappa\leq 2$, a pitchfork bifurcation occurs at $\kappa=2$, and a unique (up to rotation) nontrivial branch exists for $\kappa>2$ [9]. A significant body of work, culminating in the uniform-in-time propagation-of-chaos result of [23], now gives a remarkably complete picture of the noisy Kuramoto model.
By contrast, a quantitative description of the noisy Transformer dynamics remains largely open. Phenomena such as metastability suggest that the strong uniform-in-time results available for the noisy Kuramoto model may not hold in this richer setting; see, e.g., [28, 29]. Moreover, many natural variations, such as introducing common noise, exploring different geometries, modifying the interaction kernel, or considering anisotropic or multiplicative noise, lead to further mathematical challenges. Understanding these variants, and charting the full phase diagram of noisy attention dynamics, presents a wide landscape of open problems.
@@ Acknowledgments.
I am most grateful to Borjan Geshkovski, Shi Chen, Zhengjiang Lin, and Krishna Balasubramanian for many insightful comments on early versions of this text.